OpenStudy (aytek):

x²y''-2xy'+2y=4x3 find the general solution of this differential equation.Last of term is 4x(cube).

OpenStudy (freckles):

what ways can we use

OpenStudy (aytek):

cauchy-euler method

OpenStudy (aytek):

x power to alpha

OpenStudy (freckles):

so we can to assume the solution is \[y=x^r \\ \text{ and so the following follows } \\ y'=r x^{r-1} \\ y''=r(r-1)x^{r-2} \\ \text{ plug \in } \\ x^2 \cdot r (r-1)x^{r-2}-2x \cdot r x^{r-1}+2 \cdot x^{r}=4x^3\]

OpenStudy (freckles):

oops that 4x^3 was suppose to be 0

OpenStudy (freckles):

\[y=x^r \\ \text{ and so the following follows } \\ y'=r x^{r-1} \\ y''=r(r-1)x^{r-2} \\ \text{ plug \in } \\ x^2 \cdot r (r-1)x^{r-2}-2x \cdot r x^{r-1}+2 \cdot x^{r}=0\] solve for r

OpenStudy (aytek):

ok i solved it with cauchy euler method

OpenStudy (freckles):

so you are saying you already got the homoegenous solution'?

OpenStudy (aytek):

yes and i also found partial solution

OpenStudy (freckles):

can we use reduction of order after the euler thing?

OpenStudy (aytek):

i do not know but with cauchy it will be found the solution

OpenStudy (anonymous):

\[x^2y''-2xy'+2y=4x^3\] Let \(x=e^t\), or \(t=\ln x\), so that \(y(x)=y(e^t)=z(t)\). This means \[y'=\frac{\mathrm{d}y}{\mathrm{d}x}=\frac{1}{x}\frac{\mathrm{d}z}{\mathrm{d}t}=\frac{z'}{x}\\ y''=\frac{\mathrm{d}^2y}{\mathrm{d}x^2}=\frac{1}{x^2}\left(\frac{\mathrm{d}^2z}{\mathrm{d}t^2}-\frac{\mathrm{d}z}{\mathrm{d}t}\right)=\frac{z''-z'}{x^2}\] Substituting into the original ODE, you have \[(z''-z')-2z'+2z=4e^{3t}\implies z''-3z'+2z=4e^{3t}\] You seem to already have the homogeneous solution worked out, but I'll include it here just for the sake of completeness. The characteristic solution and its roots are \[r^2-3r+2=(r-2)(r-1)=0\implies r_1=1\quad\text{and}\quad r_2=2\]so you know that two solutions are \(C_1e^t\) and \(C_2e^{2t}\). For the nonhomogeneous part, you can try a particular solution of the form \(z_p=Ae^{3t}\), which has derivatives \[{z_p}'=3Ae^{3t}\quad\text{and}\quad{z_p}''=9Ae^{3t}\] Substituting into the ODE, you get \[9Ae^{3t}-3\left(3Ae^{3t}\right)+2Ae^{3t}=4e^{3t}\]which you can then solve for \(A\) and find \(z_p\). Keep in mind that the original ODE is in \(y\), not \(z\), so getting the final solution is just a matter of replacing \(t\) with \(\ln x\) (and simplifying, if you so choose).